The ratio of the number of sides of two regular polygons is 1:2, and the ratio of the sum of their interior angles is 3:4. Find the number of sides of each polygon.
Could you plese explain how to solve the problem.

Answer by scott8148(6628)

sum of interior angles is 180(n-2) where n is the number of sides (and angles)

let x="sides of first", so 2x="sides of second"

(180(x-2))/(180(2x-2))=3/4

"cross" multiplying __ 4(180(x-2))=3(180(2x-2)) __ 720x-1440=1080x-1080

-360=360x __ x=-1

there appears to be something wrong with the problem __ check the parameters